the sum of the ages of two men equals99.If the inverted age of the older is added to the age of the younger the sum is 108.However is the age of the younger is inverted and subtracted from the age of the older, the difference is 44. find the age of the older man
Answer by stanbon(75887)

the sum of the ages of two men equals99.If the inverted age of the older is added to the age of the younger the sum is 108.However is the age of the younger is inverted and subtracted from the age of the older, the difference is 44. find the age of the older man
-----
Equations:
x + y = 99
x + 1/y = 108
y - 1/x = 44
---------
Substitute for "x" and solve for "y" in x + (1/y) = 108
----
99-y + 1/y = 108
99y - y^2 + 1 = 108y
y^2 + 9y -1 = 0
---
y = [-9 +- sqrt(81-4*-1)]/2
Positive solution:
y = [-9 + sqrt(86)]/2
y = 0.1368 yrs
---
Comment:
I very much doubt that answer, but it is the answer
to the problem that you posted.
Cheers,
Stan H.